Please answer the following questions!! Thanks!

  1. You are considering the purchase of a common stock that paid a dividend of $3.00 yesterday. You expect this stock to have a growth rate of 20 percent for the next 3 years and the long-run normal growth rate after year 3 is expected to be constant at 5 percent. If you require a 14 percent rate of return, the price per share that you should you be willing to pay for this stock?
  2. The Moore Corporation has operating income (EBIT) of $2,250,000. The company's depreciation expense is $450,000. Moore is 120% equity financed, and it faces a 40% tax rate. What is the company's net income?
  3. Assuming all CDs have equal risk, which of the following CD's investments has the highest effective annual return (EAR)? A) A bank CD that pays 8.78 percent compounded daily. B) A bank CD that pays 9.01 percent compounded monthly. C) A bank CD that pays 9.10 percent compounded quarterly. D) A bank CD that pays 9.17 percent compounded semiannually.
  4. Which one of the following is a source of cash? a) Increase in accounts receivable, b) Decrease in accrued liabilities, c) Decrease in bank loan, d) Increase in retained earnings
Answer #1

If you require a 14 percent rate of return, the price per share that you should you be willing to pay for this stock?
=3*(1.2/1.14)+3*(1.2/1.14)^2+3*(1.2/1.14)^3+3*(1.2/1.14)^3*1.05/(14%-5%)=50.8033241

What is the company's net income?
=(2250000-450000)*(1-40%)=1080000

Assuming all CDs have equal risk, which of the following CD's investments has the highest effective annual return (EAR)?
C) A bank CD that pays 9.10 percent compounded quarterly.

Which one of the following is a source of cash?
d) Increase in retained earnings